# Rapid and Safe Trajectory Planning over Diverse Scenes through Diffusion Composition

### v2

@@ I Introduction
Traditional trajectory planners for complex environments typically comprise two phases: initial path generation and subsequent trajectory refinement via optimization. While search- and sampling-based methods can produce initial paths [1], their runtime grows rapidly with scene complexity [2], and the resulting paths are often not kinematically feasible [3]. The subsequent optimization step can enforce feasibility and safety, but adds computational burden. For time-critical applications such as planning for autonomous vehicles, the key challenge is to jointly achieve low computational cost, kinematic feasibility, and reliable collision avoidance.
Imitation learning offers an effective alternative to classical two-stage planners by directly learning from optimization-generated trajectories and modeling the distribution of near-optimal solutions [4], thereby avoiding repetitive global search and post hoc optimization. Among imitation learning methods, the diffusion model is widely used due to its ease of training [5], robust generation sampling process [6], and natural support for multi-modal inputs [7]. However, the diffusion model suffers from slow trajectory generation due to its iterative denoising process, which can lead to failure in real-world deployment. In addition, such imitation learning methods heavily rely on the training data – datasets of suboptimal kinematics or limited diversity can easily compromise performance.
@@
To harden deployment, we propose a lightweight safety filter to further enhance planning safety in real-world applications. We validate the proposed pipeline on the popular F1TENTH platform [8], which is an efficient testbed for planning methods that integrates perception, planning, and control in a unified hardware system. Experiments on a custom F1TENTH vehicle confirm the superior real-world performance of our framework.
@@ II Preliminaries
In this section, the method of path planning using potential fields [9] and its drawbacks are first introduced. Then, the approach of using Energy-based Models (EBMs) for potential field modeling is discussed in  [10]. Finally, the relationship between diffusion models and EBMs is explored.
@@ II-A Path Planning using Potential Field
Collision-free path planning can be achieved using the potential-based approach [10]. In this method, a potential function $\mathcal{U}(\bm{\tau}):\mathbb{R}^{n\times L}\to\mathbb{R}$ is used to assign a potential value to a path $\bm{\tau}$, based on predefined attractive and repulsive potentials. Here, $n$ represents the state dimension and $L$ is the path horizon. Given an initial path $\bm{\tau}$, the collision-free path can be derived by applying gradient descent to the potential function as follows: $\bm{\tau}^{k}_{1:L}=\bm{\tau}^{k-1}_{1:L}-\lambda\nabla_{\bm{\tau}_{1:L}^{k}}\mathcal{U}(\bm{\tau}_{1:L}^{k}),$ where $\bm{\tau}^{k}_{1:L}$ is the path at step $k$, and $\bm{\tau}^{k-1}_{1:L}$ is the path at the previous step, and $\lambda$ is the step size for an iteration. However, this method leads to corner cases, making it unsuitable for complex scenarios.
@@ II-B Energy-based Models
Energy-based Models are a class of generative models [11]. They can serve as surrogate models for the potential function, represented by $\mathcal{U}(\tau)\sim E_{\theta}(\bm{\tau})$, addressing the limitations of conventional potential-based methods. Given a path variable $\bm{\tau}$, the probability density $p_{\theta}(\bm{\tau})$ of a collision-free path $\bm{\tau}$ is defined as: $p_{\theta}(\bm{\tau})\propto e^{-E_{\theta}(\bm{\tau})}$, where $E_{\theta}(\bm{\tau}):\mathbb{R}^{n\times L}\to\mathbb{R}$ is an energy-based model, typically implemented using a U-Net architecture. Thus, the path planning process can be modified as follows: $\bm{\tau}_{1:L}^{t-1}=\bm{\tau}_{1:L}^{t}-\lambda\nabla_{\bm{\tau}_{1:L}^{t}}E_{\theta}(\bm{\tau}_{1:L}^{t},t,\mathcal{C})+\eta,$ where $\eta\sim\mathcal{N}(\mathbf{0},{\sigma}_{t}^{2}\bm{I})$, with $\lambda$ and $\sigma_{t}$ being specific scaling values. $t$ is the denoising step. And $\mathcal{C}=\{\mathbf{s}_{s},\mathbf{s}_{g},\mathcal{O}\}$ is the set of constraints, which includes the start state and goal state constraints denoted as $\mathbf{s}_{s}$ and $\mathbf{s}_{g}$, as well as the obstacle pose constraints $\mathcal{O}$. In this paper, $\tau=[\mathbf{s}_{1},\cdots,\mathbf{s}_{L}]_{4\times L}$ denotes a discrete planned path. Specifically, $\mathbf{s}=[x,\,y,\,q_{z},\,q_{w}]^{\top}$ with $x,\,y]^{\top}$ representing the Cartesian position and $q_{z},\,q_{w}]^{\top}$ the components of the quaternion encoding the yaw angle.
@@ III MPC-based Data Collection
For collision avoidance in the static scene, global collision-free paths are generated using the iterative collision avoidance (ItCA) method [3]. The ItCA path planning accounts for the smoothness of gear shifting points (GSPs), ensuring the kinematic feasibility of planned paths. Then, an MPC planner is employed for path tracking to produce the final global trajectory. The entire trajectory enriches the static dataset, implicitly encoding velocity through the density of trajectory points within a fixed interval $\Delta t$. Implicit learning of velocity enables more complex trajectories, exhibiting behaviors such as decelerating to yield or accelerating to bypass.
The vehicle model $f(\cdot)$ used for MPC is the kinematic model employed in [2]. The state is represented as $\bm{\zeta}=[x,\,y,\,\varphi,\,v]^{\top}$ and the control input as $\mathbf{u}=[\delta,\,a]^{\top}$. $\delta$ is the steering angle, and $a$ represents acceleration. $v$ is the velocity and $\varphi$ represents the vehicle heading.
@@
$\min_{\bm{\zeta},\mathbf{u}}$ $\gamma\cdot J_{\textrm{dynamic}}+(1-\gamma)\cdot\Big(\sum_{k=1}^{N_{p}}\overbrace{\left(\|\mathbf{A}\cdot\bm{\zeta}_{k}-\mathbf{p}_{s_{k}}^{\text{ref}}\|^{2}_{Q_{1}}\right)}^{J_{\textrm{static}}}$ (1)
$+\sum_{k=1}^{N_{p}-1}\underbrace{\|\Delta\mathbf{u}_{k}\|^{2}_{R_{1}}+\|\mathbf{A}\cdot\bm{\zeta}_{N_{p}}-\mathbf{p}_{s_{N_{p}}}^{\text{ref}}\|^{2}_{Q_{2}}+\|\mathbf{u}_{k}\|^{2}_{R_{2}}\Big)}_{J_{\textrm{static}}}$
@@
where $\mathbf{A}=[1,\,1,\,1,\,0]^{\top}$, and $N_{p}$ denotes the prediction horizon. $\bm{\zeta}_{\textrm{cur}}$ is the current vehicle state. In dynamic scenes, $\gamma=1$; otherwise, $\gamma=0$. Following the concept of projected velocity in [12, 13], the value of $s_{k}$ is determined by a fixed projection velocity $v_{\text{ref}}$, such that $s_{k}=s_{k-1}+\Delta t\cdot v_{\text{ref}}$, with $s_{0}=s_{\text{cur}}$. $\bm{\zeta}_{\text{max}}$ and $\bm{\zeta}_{\text{min}}$ denote the upper and lower bounds of the state constraints, while $\mathbf{u}_{\text{max}}$ and $\mathbf{u}_{\text{min}}$ are the upper and lower limits of the control inputs, respectively. $T_{s}$ represents the prediction interval. $Q_{1}$, $Q_{2}$, $R_{1}$, and $R_{2}$ are coefficient matrices.
@@ IV The Proposed Rapid and Safe Trajectory Planning Method
Figure 3: The Overall Framework of the Proposed Rapid and Safe Trajectory Planning (RSTP) Method. Offline Dataset Generation (Left): the ItCA [3] and MPC-based methods provide kinematically feasible trajectory datasets for training. Diffusion Composition (Middle): individual diffusion models can be flexibly composed to tackle novel scenarios not covered in the training data. Online Inference and Control (Right): The ego vehicle performs real-time inference from its current pose, and the safety filter selects the optimal trajectory per cycle for tracking.
@@ IV-A Diffusion Models Composition
where $N_{\text{all}}$ is the number of models to be composed and $\mathcal{O}_{i}\in\mathcal{O}_{\text{all}}$. The obstacles $\mathcal{O}_{i}$ are encoded using a Transformer architecture as adapted from [10]. The probability density of a collision-free trajectory is referred to as the unconditional probability density $p^{\text{traj}}_{\theta_{\text{uncond}}}(\bm{\tau}_{1:L}^{t})$. This distribution is directly learned from the dataset without explicit obstacle constraints. Each $p^{\text{traj}}_{\theta}$ corresponds to a diffusion model. And the $p^{\text{traj}}_{\theta_{\text{uncond}}}(\bm{\tau}_{1:L}^{t})$ and conditional diffusion model $p^{\text{traj}}_{\theta}(\bm{\tau}_{1:L}^{t}\mid\mathcal{O}_{i})$ can be obtained by masking the obstacle-encoder input with $\mathbf{K}=[1,0]^{\top}\otimes\mathcal{O}_{i}$ during inference, allowing the same model to output both conditional and unconditional results.
@@ V-A Environments and Baselines
Details of the experimental setup are provided in Appendix A. The diffusion model adopts the architecture and hyperparameters of [10], and its composition formula follows (3). In all experiments, the weights of the terms in Eq.(4) are set equal. The following scenes and experimental setups are established for simulation and real-world testing:
@@
We compare our RSTP method with multiple existing approaches. Baseline methods for comparison include $\text{A}^{\star}$[14], $\text{RRT}^{\star}$[15], and $\text{FMT}^{\star}$[16], all of which disregard vehicle kinematics. Additionally, kinematics-aware methods, Kino RRT[17] and $\text{Hybrid A}^{\star}$[1] are included for comparison. We denote the baseline method $\text{Hybrid A}^{\star}$ as BLM. The optimization-based methods evaluated are Rapid Iterative Trajectory Planning (RITP) [3] and Optimization-Based Collision Avoidance (OBCA) [2], both of which are initialized using paths generated by $\text{Hybrid A}^{\star}$. To assess kinematic feasibility, the MPC tracker described in Section IV is employed as the tracking controller. The mean tracking error (M.TE) serves as the metric for evaluating kinematic feasibility. Computational efficiency is evaluated by measuring the computation time (C.T). For the diffusion model, the C.T is the time to generate the final collision-free trajectory, including replanning. The percentage reduction in C.T compared to the BLM is denoted as $\text{M.RP}=\frac{\text{C.T}_{\text{BLM}}-\text{C.T}}{\text{C.T}_{\text{BLM}}}$. Similarly, the percentage reduction in M.TE is represented by $\text{M.RP}=\frac{\text{M.TE}_{\text{BLM}}-\text{M.TE}}{\text{M.TE}_{\text{BLM}}}$. The configuration with the safety filter disabled is indicated as RSTP-NSF.
@@ V-C Evaluate Compositional Performance in Real-world
The real-world experiments are conducted using a self-built F1TENTH vehicle. For mapping and localization, the vehicle utilizes the Cartographer ROS framework [18]. Communication between the server and onboard computer is managed via the Message Queuing Telemetry Transport (MQTT) protocol. The F1TENTH vehicle uses the pure pursuit controller proposed in [19] to track trajectories. The RSTP method can plan sufficiently safe trajectories in the $SS$ as shown in Fig. 4.

### v3

@@ I Introduction
Traditional trajectory planners for complex environments typically comprise two phases: initial path generation and subsequent trajectory refinement via optimization. While search- and sampling-based methods can produce initial paths [dolgov2008practical], their runtime grows rapidly with scene complexity [zhang2020optimization], and the resulting paths are often not kinematically feasible [liRapidIterativeTrajectory]. The subsequent optimization step can enforce feasibility and safety, but adds computational burden. For time-critical applications such as planning for autonomous vehicles, the key challenge is to jointly achieve low computational cost, kinematic feasibility, and reliable collision avoidance.
Imitation learning offers an effective alternative to classical two-stage planners by directly learning from optimization-generated trajectories and modeling the distribution of near-optimal solutions [urain2024deep], thereby avoiding repetitive global search and post hoc optimization. Among imitation learning methods, the diffusion model is widely used due to its ease of training [janner2022planning], robust generation sampling process [ho2020denoising], and natural support for multi-modal inputs [chi2023diffusion]. However, the diffusion model suffers from slow trajectory generation due to its iterative denoising process, which can lead to failure in real-world deployment. In addition, such imitation learning methods heavily rely on the training data – datasets of suboptimal kinematics or limited diversity can easily compromise performance.
@@
To harden deployment, we propose a lightweight safety filter to further enhance planning safety in real-world applications. We validate the proposed pipeline on the popular F1TENTH platform [betz2022autonomous], which is an efficient testbed for planning methods that integrates perception, planning, and control in a unified hardware system. Experiments on a custom F1TENTH vehicle confirm the superior real-world performance of our framework.
@@ II Preliminaries
In this section, the method of path planning using potential fields [lin2022model] and its drawbacks are first introduced. Then, the approach of using Energy-based Models (EBMs) for potential field modeling is discussed in  [luo2024potential]. Finally, the relationship between diffusion models and EBMs is explored.
@@ II-A Path Planning using Potential Field
Collision-free path planning can be achieved using the potential-based approach [luo2024potential]. In this method, a potential function $\mathcal{U}(\bm{\tau}):\mathbb{R}^{n\times L}\to\mathbb{R}$ is used to assign a potential value to a path $\bm{\tau}$, based on predefined attractive and repulsive potentials. Here, $n$ represents the state dimension and $L$ is the path horizon. Given an initial path $\bm{\tau}$, the collision-free path can be derived by applying gradient descent to the potential function as follows: $\bm{\tau}^{k}_{1:L}=\bm{\tau}^{k-1}_{1:L}-\lambda\nabla_{\bm{\tau}_{1:L}^{k}}\mathcal{U}(\bm{\tau}_{1:L}^{k}),$ where $\bm{\tau}^{k}_{1:L}$ is the path at step $k$, and $\bm{\tau}^{k-1}_{1:L}$ is the path at the previous step, and $\lambda$ is the step size for an iteration. However, this method leads to corner cases, making it unsuitable for complex scenarios.
@@ II-B Energy-based Models
Energy-based Models are a class of generative models [du2019implicit]. They can serve as surrogate models for the potential function, represented by $\mathcal{U}(\tau)\sim E_{\theta}(\bm{\tau})$, addressing the limitations of conventional potential-based methods. Given a path variable $\bm{\tau}$, the probability density $p_{\theta}(\bm{\tau})$ of a collision-free path $\bm{\tau}$ is defined as: $p_{\theta}(\bm{\tau})\propto e^{-E_{\theta}(\bm{\tau})}$, where $E_{\theta}(\bm{\tau}):\mathbb{R}^{n\times L}\to\mathbb{R}$ is an energy-based model, typically implemented using a U-Net architecture. Thus, the path planning process can be modified as follows: $\bm{\tau}_{1:L}^{t-1}=\bm{\tau}_{1:L}^{t}-\lambda\nabla_{\bm{\tau}_{1:L}^{t}}E_{\theta}(\bm{\tau}_{1:L}^{t},t,\mathcal{C})+\eta,$ where $\eta\sim\mathcal{N}(\mathbf{0},{\sigma}_{t}^{2}\bm{I})$, with $\lambda$ and $\sigma_{t}$ being specific scaling values. $t$ is the denoising step. And $\mathcal{C}=\{\mathbf{s}_{s},\mathbf{s}_{g},\mathcal{O}\}$ is the set of constraints, which includes the start state and goal state constraints denoted as $\mathbf{s}_{s}$ and $\mathbf{s}_{g}$, as well as the obstacle pose constraints $\mathcal{O}$. In this paper, $\tau=[\mathbf{s}_{1},\cdots,\mathbf{s}_{L}]_{4\times L}$ denotes a discrete planned path. Specifically, $\mathbf{s}=[x,\,y,\,q_{z},\,q_{w}]^{\top}$ with $x,\,y]^{\top}$ representing the Cartesian position and $q_{z},\,q_{w}]^{\top}$ the components of the quaternion encoding the yaw angle.
@@ III MPC-based Data Collection
For collision avoidance in the static scene, global collision-free paths are generated using the iterative collision avoidance (ItCA) method [liRapidIterativeTrajectory]. The ItCA path planning accounts for the smoothness of gear shifting points (GSPs), ensuring the kinematic feasibility of planned paths. Then, an MPC planner is employed for path tracking to produce the final global trajectory. The entire trajectory enriches the static dataset, implicitly encoding velocity through the density of trajectory points within a fixed interval $\Delta t$. Implicit learning of velocity enables more complex trajectories, exhibiting behaviors such as decelerating to yield or accelerating to bypass.
The vehicle model $f(\cdot)$ used for MPC is the kinematic model employed in [zhang2020optimization]. The state is represented as $\bm{\zeta}=[x,\,y,\,\varphi,\,v]^{\top}$ and the control input as $\mathbf{u}=[\delta,\,a]^{\top}$. $\delta$ is the steering angle, and $a$ represents acceleration. $v$ is the velocity and $\varphi$ represents the vehicle heading.
@@
$\min_{\bm{\zeta},\mathbf{u}}$ $\gamma\cdot J_{\textrm{dynamic}}+\sum_{k=1}^{N_{p}}\overbrace{\left(\|\mathbf{A}\cdot\bm{\zeta}_{k}-\mathbf{p}_{s_{k}}^{\text{ref}}\|^{2}_{Q_{1}}\right)}^{J_{\textrm{static}}}$ (1)
$+\sum_{k=1}^{N_{p}-1}\underbrace{\|\Delta\mathbf{u}_{k}\|^{2}_{R_{1}}+\|\mathbf{A}\cdot\bm{\zeta}_{N_{p}}-\mathbf{p}_{s_{N_{p}}}^{\text{ref}}\|^{2}_{Q_{2}}+\|\mathbf{u}_{k}\|^{2}_{R_{2}}}_{J_{\textrm{static}}}$
@@
where $\mathbf{A}=[1,\,1,\,1,\,0]^{\top}$, and $N_{p}$ denotes the prediction horizon. $\bm{\zeta}_{\textrm{cur}}$ is the current vehicle state. In dynamic scenes, $\gamma=1$; otherwise, $\gamma=0$. Following the concept of projected velocity in [li2024data, li2025reduce], the value of $s_{k}$ is determined by a fixed projection velocity $v_{\text{ref}}$, such that $s_{k}=s_{k-1}+\Delta t\cdot v_{\text{ref}}$, with $s_{0}=s_{\text{cur}}$. $\bm{\zeta}_{\text{max}}$ and $\bm{\zeta}_{\text{min}}$ denote the upper and lower bounds of the state constraints, while $\mathbf{u}_{\text{max}}$ and $\mathbf{u}_{\text{min}}$ are the upper and lower limits of the control inputs, respectively. $T_{s}$ represents the prediction interval. $Q_{1}$, $Q_{2}$, $R_{1}$, and $R_{2}$ are coefficient matrices.
@@ IV The Proposed Rapid and Safe Trajectory Planning Method
Figure 3: The Overall Framework of the Proposed Rapid and Safe Trajectory Planning (RSTP) Method. Offline Dataset Generation (Left): the ItCA [liRapidIterativeTrajectory] and MPC-based methods provide kinematically feasible trajectory datasets for training. Diffusion Composition (Middle): individual diffusion models can be flexibly composed to tackle novel scenarios not covered in the training data. Online Inference and Control (Right): The ego vehicle performs real-time inference from its current pose, and the safety filter selects the optimal trajectory per cycle for tracking.
@@ IV-A Diffusion Models Composition
where $N_{\text{all}}$ is the number of models to be composed and $\mathcal{O}_{i}\in\mathcal{O}_{\text{all}}$. The obstacles $\mathcal{O}_{i}$ are encoded using a Transformer architecture as adapted from [luo2024potential]. The probability density of a collision-free trajectory is referred to as the unconditional probability density $p^{\text{traj}}_{\theta_{\text{uncond}}}(\bm{\tau}_{1:L}^{t})$. This distribution is directly learned from the dataset without explicit obstacle constraints. Each $p^{\text{traj}}_{\theta}$ corresponds to a diffusion model. And the $p^{\text{traj}}_{\theta_{\text{uncond}}}(\bm{\tau}_{1:L}^{t})$ and conditional diffusion model $p^{\text{traj}}_{\theta}(\bm{\tau}_{1:L}^{t}\mid\mathcal{O}_{i})$ can be obtained by masking the obstacle-encoder input with $\mathbf{K}=[1,0]^{\top}\otimes\mathcal{O}_{i}$ during inference, allowing the same model to output both conditional and unconditional results.
@@ V-A Environments and Baselines
Details of the experimental setup are provided in Appendix A. The diffusion model adopts the architecture and hyperparameters of [luo2024potential], and its composition formula follows (3). In all experiments, the weights of the terms in Eq.(4) are set equal. The following scenes and experimental setups are established for simulation and real-world testing:
@@
We compare our RSTP method with multiple existing approaches. Baseline methods for comparison include $\text{A}^{\star}$[liu2017global], $\text{RRT}^{\star}$[karaman2011sampling], and $\text{FMT}^{\star}$[wu2021st], all of which disregard vehicle kinematics. Additionally, kinematics-aware methods, Kino RRT[sakcak2019sampling] and $\text{Hybrid A}^{\star}$[dolgov2008practical] are included for comparison. We denote the baseline method $\text{Hybrid A}^{\star}$ as BLM. The optimization-based methods evaluated are Rapid Iterative Trajectory Planning (RITP) [liRapidIterativeTrajectory] and Optimization-Based Collision Avoidance (OBCA) [zhang2020optimization], both of which are initialized using paths generated by $\text{Hybrid A}^{\star}$. To assess kinematic feasibility, the MPC tracker described in Section IV is employed as the tracking controller. The mean tracking error (M.TE) serves as the metric for evaluating kinematic feasibility. Computational efficiency is evaluated by measuring the computation time (C.T). For the diffusion model, the C.T is the time to generate the final collision-free trajectory, including replanning. The percentage reduction in C.T compared to the BLM is denoted as $\text{M.RP}=\frac{\text{C.T}_{\text{BLM}}-\text{C.T}}{\text{C.T}_{\text{BLM}}}$. Similarly, the percentage reduction in M.TE is represented by $\text{M.RP}=\frac{\text{M.TE}_{\text{BLM}}-\text{M.TE}}{\text{M.TE}_{\text{BLM}}}$. The configuration with the safety filter disabled is indicated as RSTP-NSF.
@@ V-C Evaluate Compositional Performance in Real-world
The real-world experiments are conducted using a self-built F1TENTH vehicle. For mapping and localization, the vehicle utilizes the Cartographer ROS framework [hess2016real]. Communication between the server and onboard computer is managed via the Message Queuing Telemetry Transport (MQTT) protocol. The F1TENTH vehicle uses the pure pursuit controller proposed in [weng2024aggressive] to track trajectories. The RSTP method can plan sufficiently safe trajectories in the $SS$ as shown in Fig. 4.